Please help me ASAP(for tom)
19. the tens digit of a number is six more than the ones digit. write the algebraic expression for the number.
20. the units digit of a number is four less than the tens digit. (algebraic expression)
21. the length of a rectangle is half of 5 times the width. write the expression for the perimeter of the rectangle
22. the longer base of a trapezoid is 3 morethan twice the shorter base. write an expressio for the sum of the bases.
THANKS

Found 2 solutions by rothauserc, stanbon:
Answer by rothauserc(4718)

19) let x be the ones digit and y be the tens digit, then
y = x +6
20) let x be the tens digit and y be the ones digit, then
y = x -4
21) let l be the length and w be the width, then
l = (1/2)*5w
Perimeter(P) = 2*((1/2*5w) + 2w
2P = 10w + 4w
P = 5w +2w = 7w
22) let x be the shortest base and y be the longer base, then
y = 2x + 3
sum of the bases = x + 2x +3 = 3x +3


Answer by stanbon(75887)


Let the number be 10t+u
--------------------
19. the tens digit of a number is six more than the ones digit. write the algebraic expression for the number.
t = u + 6
# = 10(u+6)+u = 11u + 60
------------------------------
20. the units digit of a number is four less than the tens digit. (algebraic expression)
u = t-4
# = 10t + t-4 = 9t-4
------------------------------
21. the length of a rectangle is half of 5 times the width. write the expression for the perimeter of the rectangleth
P = 2(length + width)
P = 2((1/2)5w+w) = 5w+2w = 7w
----------------------
22. the longer base of a trapezoid is 3 more than twice the shorter base. write an expressio for the sum of the bases.
b1 = 2*b2 + 3
Sum + b2 + 2b2+3 = 3b2+3
-------------
Cheers,
Stan H.
----------------